OpenStudy (anonymous):

14+3*5^x=389 solved for x please

OpenStudy (anonymous):

subtract both sides by 14. divide both sides by 3. now, use logs. a^b = c -> b = log a (c)

OpenStudy (anonymous):

\[a^b = c\] \[\log_{a}c = b \]

OpenStudy (anonymous):

okay so what is 5^x=372 solved for x